OpenStudy (anonymous):

simplify:((2x^2+9x+9)/(x+1))/((10x^2+19x+6)/(5x^2+7x+2))

OpenStudy (anonymous):

and yes this is a fraction within a fraction

OpenStudy (anonymous):

where is dumbcow?

OpenStudy (anonymous):

everything will be easier when you factor, in the first fraction factor the numerator.In the second fraction factor both the numerator and denominator

OpenStudy (anonymous):

times by reciprical and solv

OpenStudy (dumbcow):

yeah what they said

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

that is an insane polyomial there

OpenStudy (anonymous):

ya our school does that to us sometimes

OpenStudy (anonymous):

I know what you mean. I love it when our SI says "What he didn't teach you this" a week after class

OpenStudy (anonymous):

"Telescoping a series for divergence is not needed".... Well when the book asks you...

OpenStudy (dumbcow):

whenever you are dividing by a fraction, rewrite so you are multiplying by reciprocal \[\frac{\frac{1}{x}}{\frac{3}{x}} = \frac{1}{x}*\frac{x}{3}\]

OpenStudy (anonymous):

or finding the sum of a sequence i mena

OpenStudy (anonymous):

ok dumbcow you make it so much clearer.. thanks

OpenStudy (anonymous):

and you too outkast

OpenStudy (dumbcow):

hows the factoring going? everything factors and almost everything cancels answer -> x+3